# Extremal random matrices with independent entries and matrix superconcentration inequalities

### v1

@@ 1. Introduction
In the special case that $b_{ij}=1$ for all $i,j$, we recover one of the most classical models of random matrix theory: then $\|X\|^{2}=\|X^{*}X\|$ is the norm of a Wishart matrix with unit covariance. In this case, classical methods of random matrix theory yield the exact asymptotic behavior [11]
@@
where $F_{1}(s)$ is the distribution function of the Tracy-Widom law of order one. From this expression, we immediately read off that in the above asymptotic regime $\|X\|=(1+o(1))(\sqrt{n}+\sqrt{m})$ with fluctuations of order $n^{-1/6}\vee m^{-1/6}$. It is important to note that the scale of the fluctuations is much smaller than what is predicted by general concentration of measure principles [13], which yield fluctuations of order $O(1)$ in the present setting. The presence of such unexpectedly small fluctuations is sometimes called the superconcentration phenomenon [6].
@@
The following is shown in [3, Corollary 3.11].
Theorem 1.1 ([3]).
@@
with high probability as $n\to\infty$. It follows that the second term in (1.3) dominates the behavior of $\|X\|$ in the sparse regime $k\ll\log n$. In fact, (1.3) optimally captures the phase transition for the leading order behavior of the norm of random band matrices (cf. [3, Corollary 4.4] in the self-adjoint case).
@@ 1.1.1. Small deviations
for $0\leq t\leq n^{1/6}\sqrt{m}$. This captures precisely the fluctuations and the upper tail of the Tracy-Widom asymptotics (1.1), up to a universal constant. To the best of our knowledge, the nonasymptotic bound is new even in this very special case.
@@ 1.1.2. Large deviations
Theorem 1.4 controls small deviations, that is, deviations up to the order of the mean. In contrast, the large deviations of $\|X\|$ are controlled by its Gaussian tail behavior [17, Corollary 3.2], and we cannot expect a qualitative improvement over Theorem 1.1 in this setting.
@@ 1.1.3. Extremum principle
To explain this principle, let us begin by recalling the idea behind the proof of Theorem 1.1. Rather than bound the norm of the nonhomogeneous model $X$ directly, the approach of [3] compares $X$ with an i.i.d. matrix of smaller dimension, whose norm can be controlled by any classical method for homogeneous random matrices. In particular, the main technical device in [3] shows that when normalized so that $\sigma_{*}=1$, the $2p$-moment of $X$ can be bounded by the $2p$-moment of a matrix with i.i.d. entries of dimension $\lceil\sigma_{1}^{2}+p\rceil\times\lceil\sigma_{2}^{2}+p\rceil$. This suffices to capture the leading order behavior of $\|X\|$, but the dependence of the comparison matrix on $p$ precludes any accurate control of the fluctuations.
@@
Note that the inequality in Theorem 1.7 holds with equality when $X$ has i.i.d. entries. Thus, in contrast to the comparison principle of [3], Theorem 1.7 may be viewed as an extremum principle for random matrices. In terms of normalized moments, this principle may be expressed as follows.
@@
Theorem 1.7 will be proved in section 4 below. With this result in hand, the proof of Theorems 1.4 and 1.6 will reduce to the proof of analogous tail bounds for Wishart matrices. The latter is made possible by a beautiful method, pioneered by Ledoux [14, 15, 16], for deriving Tracy-Widom type tail bounds from sharp moment estimates. While such estimates were obtained by Ledoux for GUE [14] and GOE [16] matrices, sharp moment bounds for Wishart matrices do not appear to be known in the literature. Such bounds are obtained here in section 5. The Wishart model is considerably more delicate than the GUE and GOE models due to the fact that we require bounds that hold uniformly for all dimensions $n,m$, not just in the asymptotic regime (1.1) where $m=cn$ are proportional.
Let us emphasize that, despite the apparent similarity between Theorem 1.7 and the comparison principles of [3], the proofs of these results are completely different. The method of [3] uses a classical combinatorial expression for the moments as a sum over equivalence classes of even closed walks, and estimates each term separately. In contrast, the proof of Theorem 1.7 expresses the moments of a Gaussian random matrix as a sum over pairings by the Wick formula, and then uses an iterative procedure to reduce each term to a noncrossing pairing. This new approach turns out to provide a much more efficient mechanism for controlling the moments.
@@ 2. $\bullet$
    We then consider in section 3 self-adjoint random matrices with independent real entries, that is, the self-adjoint analogue of Model 1.3. The proof of the the extremum principle is much more involved in this case, and shares the same difficulties as the proof of the extremum principle for non-self-adjoint models. However, for both complex and real self-adjoint models, we can directly invoke the moment estimates of Ledoux [14, 15, 16] to obtain tail bounds.
@@ 1.2.1. Intrinsic freeness
    The extremum principles of [3] and of this paper compare the spectral statistics of $X$ with those of an i.i.d. random matrix $Y$ of smaller dimension.
@@
    The intrinsic freeness principle of [1] compares the spectral statistics of $X$ with those of a deterministic operator $X_{\rm free}$ that arises in free probability theory.
The theory of [1] is in fact much more general than the setting of this paper, in that it captures both random matrices with dependent entries and more general spectral statistics. However, the following special case may be directly compared to the bounds of this paper: if $X$ is an $n\times m$ random matrix with independent Gaussian entries $X_{ij}\sim N(0,b_{ij}^{2})$ and $n\leq m$, then [1, Corollary 2.2] yields
@@
for every $\varepsilon>0$ (this may be shown as in [1, Lemma 3.1]). As
@@
by [1, Lemma 2.5], the intrinsic freeness principle readily yields a slightly weaker form of (1.3). Note that the second term in (1.6) is even larger than that of (1.3), while the results of this paper yield a much smaller second-order term.
@@
We can apply [1, Lemma 3.2 and Remark 2.6] to obtain
@@
as $n\to\infty$ (cf. [1, 2] and [19, §2.1]). Moreover, it is shown in [19, Theorem 3] that when $\delta>1$, the fluctuations of $\|X\|$ are of order $\sigma_{*}\sim n^{-1/2}$. This shows that we cannot replace $\sigma_{1}+\sigma_{2}$ by $\|X_{\rm free}\|$ in Theorem 1.4, as that would imply a much smaller bound on the fluctuations of order $\sigma_{*}^{4/3}$.
Example 1.11 illustrates in a particularly clear manner that the extremum and intrinsic freeness principles capture fundamentally different mathematical phenomena. This may appear somewhat surprising, given that both the proofs in this paper and those of [1] are based on the control of crossings in the Wick formula (in the latter case, this idea is due to [22]). However, these results exploit crossings in very different ways: here we show that the contribution of each crossing to the nonhomogeneous model is bounded by that in the i.i.d. model, while the intrinsic freeness principle aims to show that crossings are negligible altogether.
@@ 1.2.2. Superconcentration
The concentration of measure phenomenon [13] is a powerful tool for bounding the fluctuations of random structures. While general concentration inequalities often yield optimal bounds, there are also situations where the scale of the flucutations is much smaller than is predicted by general principles. This phenomenon was called superconcentration in [6].
When applied to the random matrix models of this paper, classical concentration inequalities ensure that the scale of the fluctuations of $\|X\|$ is at most of order $\sigma_{*}$, cf. [1, Corollary 4.14]. In contrast, the fluctuation term in Theorem 1.4, of order $\sigma_{*}^{4/3}\sigma_{1}^{-1/3}$, is often much smaller than $\sigma_{*}$. We may therefore think of Theorem 1.4 as a “matrix superconcentration inequality”. Indeed, this bound accurately captures the well known superconcentration property of the upper tail of the norm of Wishart matrices, and extends it to a much larger class of models. [^1] After the completion of this paper, we learned that a weaker superconcentration property was previously obtained by entirely different methods in [9, Theorem 3.4] for self-adjoint random matrices satisfying the analogue of Lemma 1.10. Theorem 3.4 below yields a much more precise tail bound in the self-adjoint setting that is essentially optimal by the extremum principle.
@@
More generally, it is not expected that the specific setting of Lemma 1.10 is necessary for superconcentration. General principles explained in [6] suggest that the spectral norm should exhibit superconcentration provided there are many singular values near the maximal one (this fails in Example 1.11, where the largest singular value is isolated from the bulk). Furthermore, even in the setting of Lemma 1.10, our results yield an upper bound that is sharp for block-diagonal matrices, but other such models may exhibit even smaller fluctuations (e.g., [21]). A precise understanding of when and how much superconcentration arises for nonhomogeneous random matrices remains out of reach of any known method.
@@ 1.2.3. Universality
The setting of Model 1.3 requires the entries of $X$ to be symmetrically distributed and have all their moments dominated by those of the Gaussian distribution. However, the classical Tracy-Widom asymptotics (1.1) remain valid in a much more general setting [20]: it suffices that the entries of $X$ have zero mean, unit variance, and soft control of the higher moments.
It is of considerable interest to understand whether nonasymptotic bounds can also be achieved in this much more general setting. An extension of Theorem 1.1 along these lines was proved in [12, §4.3]: it was shown there that for any random matrix $X$ whose entries are independent and have zero mean, the statement of Theorem 1.1 remains valid if we replace the parameters (1.2) by
@@
However, just as Theorem 1.1, this result fails to capture fluctuations at the Tracy-Widom scale. It is natural to conjecture the validity of an analogous extension of Theorem 1.4 and of the other main results of this paper. Such results would be of particular interest in applications, e.g., to random graphs [4].
Unfortunately, it is not clear whether the methods of this paper can be adapted to achieve such results (nor do they follow from the very general universality principles in [5], whose error terms are far larger than the fluctuations in Theorem 1.4). While much less precise, the method used in [3, 12] does not use any special properties of the Gaussian distribution and is therefore readily adapted to more general models. In contrast, the proof of the sharp extremum principle of Theorem 1.7 is based on the Wick formula, and is therefore inherently Gaussian in nature. For this reason, an analogous extension of our main results remains an open problem.
@@ 2.1. Extremum principle
by the classical Wick formula [18, Theorem 22.3 and Remark 22.5]. ∎
Free probability theory suggests [18, p. 367] that the expression in Lemma 2.5 should be dominated by noncrossing pairings. The idea behind the proof of Theorem 2.2 is to consider separately the effect of noncrossing and crossing pairs on the terms in the sum in Lemma 2.5. We first consider noncrossing pairings.
@@
with equality when $b_{ij}=1$ for all $i,j$. Here we denote !1:=*!1:=*! 1 := * and !*:=1!*:=1! * := 1.
@@
An identity of the above form for $\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]$ is classical in random matrix theory, where it is known as the genus expansion [18, Theorem 22.12]. In particular, the precise combinatorial meaning of $\ell(\pi)$ can be understood in terms of the genus of the orientable surface obtained by gluing together the edges of a regular $2p$-gon corresponding to each pair of $\pi$. However, this combinatorial intepretation is completely irrelevant for our purposes: all we used is that the inequalities we apply to the nonhomogeneous matrix $X$ become equality for $Y$. For the real random matrices investigated in the following sections, the combinatorial structure is much more delicate while a comparison argument remains tractable.
@@ 2.2. Small deviations
We can now combine Theorem 2.2 with a method of Ledoux [15, §5.2] to obtain small deviations inequalities at the Tracy-Widom scale. The basis for this method is an accurate estimate on the $p$th moment of $Y$ for moderately large $p$. The following can be read off from [15, pp. 210–211].
@@ 2.3. Large deviations
The moment estimate of Lemma 2.9 is not accurate for large $p$: indeed, this estimate yields a subexponential bound $\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]^{1/2p}=O(p)$ as $p\to\infty$, while the large deviations of the norms of Gaussian random matrices are in fact subgaussian. To obtain a large deviation inequality, we will instead employ the following estimate of Haagerup and Thorbjørnsen [10, Eq. (3.5)].
@@
This yields the following.
Proposition 2.12.
For $X$ as in Model 2.1 and $\varepsilon\geq 0$, we have
$\mathbf{P}\big{[}\|X\|>2\sqrt{\sigma^{2}+\sigma_{*}^{2}}+\sigma_{*}\varepsilon\big{]}\leq 2n\,e^{-\varepsilon^{2}/2}.$
Proof.
@@
for all $\varepsilon\geq 0$, where we used $d\leq\sigma^{2}+1$. This concludes the proof for $\sigma_{*}=1$. For general $\sigma_{*}$, it suffices to apply the above bound to the random matrix $\frac{X}{\sigma_{*}}$. ∎
Theorem 2.4 follows readily.
Proof of Theorem 2.4.
The conclusion follows immediately from Proposition 2.12 using that $\sqrt{\sigma^{2}+\sigma_{*}^{2}}\leq\sigma+\frac{\sigma_{*}^{2}}{2\sigma}\leq\sigma+\frac{\sigma_{*}}{2}$. ∎

### v2

@@ 1. Introduction
In the special case that $b_{ij}=1$ for all $i,j$, we recover one of the most classical models of random matrix theory: then $\|X\|^{2}=\|X^{*}X\|$ is the norm of a Wishart matrix with unit covariance. In this case, classical methods of random matrix theory yield the exact asymptotic behavior [Joh01]
@@
where $F_{1}(s)$ is the distribution function of the Tracy-Widom law of order one. From this expression, we immediately read off that in the above asymptotic regime $\|X\|=(1+o(1))(\sqrt{n}+\sqrt{m})$ with fluctuations of order $n^{-1/6}\vee m^{-1/6}$. It is important to note that the scale of the fluctuations is much smaller than what is predicted by general concentration of measure principles [Led01], which yield fluctuations of order $O(1)$ in the present setting. The presence of such unexpectedly small fluctuations is sometimes called the superconcentration phenomenon [Cha14].
@@
The following is shown in [BvH16, Corollary 3.11].
Theorem 1.1 ([BvH16]).
@@
with high probability as $n\to\infty$. It follows that the second term in (1.3) dominates the behavior of $\|X\|$ in the sparse regime $k\ll\log n$. In fact, (1.3) optimally captures the phase transition for the leading order behavior of the norm of random band matrices (cf. [BvH16, Corollary 4.4] in the self-adjoint case).
@@ 1.1.1. Small deviations
for $0\leq t\leq n^{1/6}\sqrt{m}$. This captures precisely the fluctuations and the upper tail of the Tracy-Widom asymptotics (1.1), up to a universal constant. To the best of our knowledge, the nonasymptotic bound is new even in this very special case (a slightly weaker bound in this setting may be found in [LR10, Theorem 2]).
@@ 1.1.2. Large deviations
Theorem 1.4 controls small deviations, that is, deviations up to the order of the mean. In contrast, the large deviations of $\|X\|$ are controlled by its Gaussian tail behavior [LT91, Corollary 3.2], and we cannot expect a qualitative improvement over Theorem 1.1 in this setting.
@@ 1.1.3. Extremum principle
To explain this principle, let us begin by recalling the idea behind the proof of Theorem 1.1. Rather than bound the norm of the nonhomogeneous model $X$ directly, the approach of [BvH16] compares $X$ with an i.i.d. matrix of smaller dimension, whose norm can be controlled by any classical method for homogeneous random matrices. In particular, the main technical device in [BvH16] shows that when normalized so that $\sigma_{*}=1$, the $2p$-moment of $X$ can be bounded by the $2p$-moment of a matrix with i.i.d. entries of dimension $\lceil\sigma_{1}^{2}+p\rceil\times\lceil\sigma_{2}^{2}+p\rceil$. This suffices to capture the leading order behavior of $\|X\|$, but the dependence of the comparison matrix on $p$ precludes any accurate control of the fluctuations.
@@
Note that the inequality in Theorem 1.7 holds with equality when $X$ has i.i.d. entries. Thus, in contrast to the comparison principle of [BvH16], Theorem 1.7 may be viewed as an extremum principle for random matrices. In terms of normalized moments, this principle may be expressed as follows.
@@
Theorem 1.7 will be proved in section 4 below. With this result in hand, the proof of Theorems 1.4 and 1.6 will reduce to the proof of analogous tail bounds for Wishart matrices. The latter is made possible by a beautiful method, pioneered by Ledoux [Led04, Led07, Led09], for deriving Tracy-Widom type tail bounds from sharp moment estimates.111 Tail bounds for Wigner and Wishart matrices have also been proved by methods that are not based on moments, such as in [Aub05, LR10]. However, as Theorem 1.7 only provides a comparison principle for moments and not for tail probabilities, the moment approach is essential here. While such estimates were obtained by Ledoux for GUE [Led04] and GOE [Led09] matrices, sharp moment bounds for Wishart matrices do not appear to be known in the literature. Such bounds are obtained here in section 5. The Wishart model is considerably more delicate than the GUE and GOE models due to the fact that we require bounds that hold uniformly for all dimensions $n,m$, not just in the asymptotic regime (1.1) where $m=cn$ are proportional.
Let us emphasize that, despite the apparent similarity between Theorem 1.7 and the comparison principles of [BvH16], the proofs of these results are completely different. The method of [BvH16] uses a classical combinatorial expression for the moments as a sum over equivalence classes of even closed walks, and estimates each term separately. In contrast, the proof of Theorem 1.7 expresses the moments of a Gaussian random matrix as a sum over pairings by the Wick formula, and then uses an iterative procedure to reduce each term to a noncrossing pairing. This new approach turns out to provide a much more efficient mechanism for controlling the moments.
@@ 2. $\bullet$
    We then consider in section 3 self-adjoint random matrices with independent real entries, that is, the self-adjoint analogue of Model 1.3. The proof of the the extremum principle is much more involved in this case, and shares the same difficulties as the proof of the extremum principle for non-self-adjoint models. However, for both complex and real self-adjoint models, we can directly invoke the moment estimates of Ledoux [Led04, Led07, Led09] to obtain tail bounds.
@@ 1.2.1. Intrinsic freeness
    The extremum principles of [BvH16] and of this paper compare the spectral statistics of $X$ with those of an i.i.d. random matrix $Y$ of smaller dimension.
@@
    The intrinsic freeness principle of [BBV23] compares the spectral statistics of $X$ with those of a deterministic operator $X_{\rm free}$ that arises in free probability theory.
The theory of [BBV23] is in fact much more general than the setting of this paper, in that it captures both random matrices with dependent entries and more general spectral statistics. However, the following special case may be directly compared to the bounds of this paper: if $X$ is an $n\times m$ random matrix with independent Gaussian entries $X_{ij}\sim N(0,b_{ij}^{2})$ and $n\leq m$, then [BBV23, Corollary 2.2] yields
@@
for every $\varepsilon>0$ (this may be shown as in [BBV23, Lemma 3.1]). As
@@
by [BBV23, Lemma 2.5], the intrinsic freeness principle readily yields a slightly weaker form of (1.3). Note that the second term in (1.6) is even larger than that of (1.3), while the results of this paper yield a much smaller second-order term.
@@
We can apply [BBV23, Lemma 3.2 and Remark 2.6] to obtain
@@
as $n\to\infty$ (cf. [BBV23, BCSV23] and [Pau07, §2.1]). Moreover, it is shown in [Pau07, Theorem 3] that when $\delta>1$, the fluctuations of $\|X\|$ are of order $\sigma_{*}\sim n^{-1/2}$. This shows that we cannot replace $\sigma_{1}+\sigma_{2}$ by $\|X_{\rm free}\|$ in Theorem 1.4, as that would imply a much smaller bound on the fluctuations of order $\sigma_{*}^{4/3}$.
Example 1.11 illustrates in a particularly clear manner that the extremum and intrinsic freeness principles capture fundamentally different mathematical phenomena. This may appear somewhat surprising, given that both the proofs in this paper and those of [BBV23] are based on the control of crossings in the Wick formula (in the latter case, this idea is due to [Tro18]). However, these results exploit crossings in very different ways: here we show that the contribution of each crossing to the nonhomogeneous model is bounded by that in the i.i.d. model, while the intrinsic freeness principle aims to show that crossings are negligible altogether.
@@ 1.2.2. Superconcentration
The concentration of measure phenomenon [Led01] is a powerful tool for bounding the fluctuations of random structures. While general concentration inequalities often yield optimal bounds, there are also situations where the scale of the flucutations is much smaller than is predicted by general principles. This phenomenon was called superconcentration in [Cha14].
When applied to the random matrix models of this paper, classical concentration inequalities ensure that the scale of the fluctuations of $\|X\|$ is at most of order $\sigma_{*}$, cf. [BBV23, Corollary 4.14]. In contrast, the fluctuation term in Theorem 1.4, of order $\sigma_{*}^{4/3}\sigma_{1}^{-1/3}$, is often much smaller than $\sigma_{*}$. We may therefore think of Theorem 1.4 as a “matrix superconcentration inequality”. Indeed, this bound accurately captures the well known superconcentration property of the upper tail of the norm of Wishart matrices, and extends it to a much larger class of models. [^2] After the completion of this paper, we learned that a weaker superconcentration property was previously obtained by entirely different methods in [EK11, Theorem 3.4] for self-adjoint random matrices satisfying the analogue of Lemma 1.10. Theorem 3.4 below yields a much more precise tail bound in the self-adjoint setting that is essentially optimal by the extremum principle.
@@
More generally, it is not expected that the specific setting of Lemma 1.10 is necessary for superconcentration. General principles explained in [Cha14] suggest that the spectral norm should exhibit superconcentration provided there are many singular values near the maximal one (this fails in Example 1.11, where the largest singular value is isolated from the bulk). Furthermore, even in the setting of Lemma 1.10, our results yield an upper bound that is sharp for block-diagonal matrices, but other such models may exhibit even smaller fluctuations (e.g., [Sod10]). A precise understanding of when and how much superconcentration arises for nonhomogeneous random matrices remains out of reach of any known method.
@@ 1.2.3. Universality
The setting of Model 1.3 requires the entries of $X$ to be symmetrically distributed and have all their moments dominated by those of the Gaussian distribution. However, the classical Tracy-Widom asymptotics (1.1) remain valid in a much more general setting [PY14]: it suffices that the entries of $X$ have zero mean, unit variance, and soft control of the higher moments.
It is of considerable interest to understand whether nonasymptotic bounds can also be achieved in this much more general setting. An extension of Theorem 1.1 along these lines was proved in [LVY18, §4.3]: it was shown there that for any random matrix $X$ whose entries are independent and have zero mean, the statement of Theorem 1.1 remains valid if we replace the parameters (1.2) by
@@
However, just as Theorem 1.1, this result fails to capture fluctuations at the Tracy-Widom scale. It is natural to conjecture the validity of an analogous extension of Theorem 1.4 and of the other main results of this paper. Such results would be of particular interest in applications, e.g., to random graphs [BBK20].
Unfortunately, it is not clear whether the methods of this paper can be adapted to achieve such results (nor do they follow from the very general universality principles in [BV23], whose error terms are far larger than the fluctuations in Theorem 1.4). While much less precise, the method used in [BvH16, LVY18] does not use any special properties of the Gaussian distribution and is therefore readily adapted to more general models. In contrast, the proof of the sharp extremum principle of Theorem 1.7 is based on the Wick formula, and is therefore inherently Gaussian in nature. For this reason, an analogous extension of our main results remains an open problem.
@@ 2.1. Extremum principle
by the classical Wick formula [NS06, Theorem 22.3 and Remark 22.5]. ∎
Free probability theory suggests [NS06, p. 367] that the expression in Lemma 2.5 should be dominated by noncrossing pairings. The idea behind the proof of Theorem 2.2 is to consider separately the effect of noncrossing and crossing pairs on the terms in the sum in Lemma 2.5. We first consider noncrossing pairings.
@@
with equality when $b_{ij}=1$ for all $i,j$. Here we denote !1:=!1:=*! 1 := ∗ and !:=1!*:=1! ∗ := 1.
@@
An identity of the above form for $\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]$ is classical in random matrix theory, where it is known as the genus expansion [NS06, Theorem 22.12]. In particular, the precise combinatorial meaning of $\ell(\pi)$ can be understood in terms of the genus of the orientable surface obtained by gluing together the edges of a regular $2p$-gon corresponding to each pair of $\pi$. However, this combinatorial intepretation is completely irrelevant for our purposes: all we used is that the inequalities we apply to the nonhomogeneous matrix $X$ become equality for $Y$. For the real random matrices investigated in the following sections, the combinatorial structure is much more delicate while a comparison argument remains tractable.
@@ 2.2. Small deviations
We can now combine Theorem 2.2 with a method of Ledoux [Led07, §5.2] to obtain small deviations inequalities at the Tracy-Widom scale. The basis for this method is an accurate estimate on the $p$th moment of $Y$ for moderately large $p$. The following can be read off from [Led07, pp. 210–211].
@@ 2.3. Large deviations
The moment estimate of Lemma 2.9 is not accurate for large $p$: indeed, this estimate yields a subexponential bound $\mathbf{E}[\mathop{\mathrm{tr}}Y^{2p}]^{1/2p}=O(p)$ as $p\to\infty$, while the large deviations of the norms of Gaussian random matrices are in fact subgaussian. To obtain a large deviation inequality, we will instead employ the following estimate of Haagerup and Thorbjørnsen [HT03, Eq. (3.5)].
@@
We can now prove Theorem 2.4.
Proof of Theorem 2.4.
@@
for all $\varepsilon\geq 0$, where we used $d\leq\sigma^{2}+1$.
For general $\sigma_{*}$, applying the above bound to the random matrix $\frac{X}{\sigma_{*}}$ yields
$\mathbf{P}\big{[}\|X\|>2\sqrt{\sigma^{2}+\sigma_{*}^{2}}+\sigma_{*}\varepsilon\big{]}\leq 2n\,e^{-\varepsilon^{2}/2}$
for all $\varepsilon\geq 0$. To conclude, it remains to use that $\sqrt{\sigma^{2}+\sigma_{*}^{2}}\leq\sigma+\frac{\sigma_{*}^{2}}{2\sigma}\leq\sigma+\frac{\sigma_{*}}{2}$. ∎